# Ravel: Evaluating Interpretability Methods on Disentangling Language Model Representations

### v1

@@ Abstract
Mor Geva Tel Aviv University morgeva@tauex.tau.ac.il                        Atticus Geiger Pr(Ai)${}^{2}$ Group atticusg@gmail.com
@@ 1 Introduction
The metric we use to assess interpretability methods is based on interchange interventions (also known as activation patching). This operation has emerged as a workhorse in interpretability, with a wide swath of research applying the technique to test if a high-level concept is stored in a model representation and used during inference Geiger et al. (2020); Vig et al. (2020); Geiger et al. (2021); Finlayson et al. (2021); Meng et al. (2022); Chan et al. (2022); Geva et al. (2023); Wang et al. (2023); Hanna et al. (2023); Conmy et al. (2023); Goldowsky-Dill et al. (2023); Hase et al. (2023); Todd et al. (2024); Feng and Steinhardt (2024); Cunningham et al. (2024); Huang et al. (2023); Tigges et al. (2023); Lieberum et al. (2023); Davies et al. (2023); Hendel et al. (2023); Ghandeharioun et al. (2024).
Specifically, we use the LM to process a prompt like “Paris is in the continent of” and then intervene on the neurons $\mathbf{N}$ to fix the feature $F$ to be the value it would have if the LM were given a prompt like “Tokyo is a large city.” If this leads the LM to output “Asia” instead of “Europe”, then we have evidence that the feature $F$ encodes the attribute “continent”. Then, we perform the same intervention when the LM processes a prompt like “People in Paris speak”. If the LM still outputs “French”, then we have evidence that the feature $F$ has disentangled the attributes “continent” and “language”.
@@ 2 The Ravel Dataset
The goal of Ravel is to assess the ability of methods to isolate individual explanatory factors in model representations (desideratum 4), and do so in a way that is faithful to how the target models work (desideratum 1). The dataset train/test structure seeks to ensure that methods are evaluated by how well their explanations generalize to new cases (desideratum 2), and Ravel is designed to support intervention-based metrics that assess the extent to which methods have found representations that causally affect the model behavior (desideratum 3).
Entity Type
Attributes
# Entities # Prompt Templates
City
Country, Language, Latitude, Longitude, Timezone, Continent
3552 150
Nobel Laureate
Award Year, Birth Year, Country of Birth, Field, Gender
928 100
Verb
Definition, Past Tense, Pronunciation, Singular
986 60
Physical Object
Biological Category, Color, Size, Texture
563 60
Occupation
Duty, Gender Bias, Industry, Work Location
799 50
@@ 2.1 Data Generation
Ravel offers two settings, Entity and Context, to evaluate the generalizability (desideratum 2) of an interpretability method across unseen entities and contexts. Each setting has a predefined train/dev/test structure. In Entity, for each entity type, we randomly split the entities into 50%/25%/25% for train/dev/test, but use the same set of prompt templates across the three splits. In Context, for each attribute, we randomly split the prompt templates into 50%/25%/25%, but use the same set of entities across the three splits.
@@ 2.2 Interpretability Evaluation
A central goal of Ravel is to assess methods by the extent to which they provide causal explanations of model behaviors (desideratum 3). To build such analyses, we need to put models into counterfactual states that allow us to isolate the causal effects of interest.
@@
Geiger et al. (2023b) generalize this operation to intervene upon features that are distributed across neurons using a bijective featurizer $\mathcal{F}$. Let $\mathcal{M}_{F\leftarrow f}$ be the model where neurons $\mathbf{N}$ are projected into a feature space, the feature $F$ is fixed to take on value $f$, and then the features are projected back into the space of neural activations using $\mathcal{F}^{-1}$. If we let $\tau(\mathcal{M}(x))$ be the token that a model predicts for a given prompt $x\in\mathcal{D}{}{}$, then comparisons between $\tau(\mathcal{M}(x))$ and $\tau(\mathcal{M}_{F\leftarrow f}(x))$ yield insights into the causal role that $F$ plays in model behavior.
@@
If $F$ encodes $A$, then interventions on $F$ should change the value of $A$. When $\mathcal{M}$ is given a prompt $x\in\mathcal{P}_{E}^{A}$, we can intervene on $F$ to set the value to what it would be if a second prompt $x^{\prime}\in\mathcal{P}_{E^{\prime}}^{A^{\prime}}\cup\mathcal{W}_{E^{\prime}}$ was provided. The token predicted by $\mathcal{M}$ should change from $A_{E}$ to $A_{E^{\prime}}$:
@@
In practice, two attributes might not be fully disentanglable in the model $\mathcal{M}$ so there is no guarantee that it is possible to find a feature $F$ that achieves $\texttt{Cause}=1$ and $\texttt{Iso}=1$ at the same time. However, evidence that two attributes might not be separable is an insight into how knowledge is structured in the model. To balance these two objectives, we define the Disentangle score as a weighted average between Cause and Iso.
@@ 3.2 Sparse Autoencoder
$\sum_{x\in\mathcal{D}}||\texttt{GetVals}(\mathcal{M}(x),\mathbf{N})-\Big{(}W_{2}\mathbf{f}+b_{2}\Big{)}||+||\mathbf{f}||_{1}$
$\mathbf{f}=\mathsf{ReLU}(W_{1}(\texttt{GetVals}(\mathcal{M}(x),\mathbf{N})-b_{0})+b_{1})$
with $W_{1}\in\mathbb{R}^{n\times m}$, $W_{2}\in\mathbb{R}^{m\times n}$, $b_{0},b_{2}\in\mathbb{R}^{n}$, and $b_{1}\in\mathbb{R}^{m}$. The trained autoencoder is then used to define a featurizer
$\mathcal{F}_{A}(\mathbf{n})=\mathsf{ReLU}(W_{1}(\mathbf{n}+b_{0})+b_{1})$
@@ 3.3 Relaxed Linear Adversarial Probe
We define the $\mathcal{F}$ using the set of $k$ orthonormal vectors that span the row space and the set of $n-k$ orthonormal vectors that span the null space of $W$:
@@
Our feature $F_{A}$ is the first $k$ dimensions of the feature space, or simply $W$. Intuitively, since the linear probe was trained to extract the attribute $A$, the rowspace is the linear subspace of neural activations that the probe is “looking at” to make predictions.
@@ 3.5 Distributed Alignment Search
$\mathcal{L}_{\texttt{Cause}}(A,F_{A},\mathcal{M})=\texttt{CE}(\texttt{II}(\mathcal{M}(x),F_{A},x^{\prime}),A_{E^{\prime}})$
@@ 4.2 Results
We report the disentanglement score on test splits in Table 2. A breakdown of the results with per-attribute Cause and Iso is in Appendix C.
@@
MDAS is the state-of-the-art method on Ravel, being able to achieve high Disentangle scores while only intervening on a feature $F_{A}$ with a dimensionality that is $4\%\cdot|\mathbf{N}|$ where $\mathbf{N}$ are the neurons the feature is distributed across (Figure 2). DBM, MDBM, and DAS, the other methods that are trained with interventions using counterfactual labels as supervision, achieve the next best performance. PCA and Sparse Autoencoder achieve the lowest Disentangle scores, which aligns with the prior finding that disentangled representations are difficult to learn without supervision Locatello et al. (2018). Unsurprisingly, more supervision results in higher performance.
@@ 5 Related Work
Intervention-based techniques, branching off from interchange intervention (Vig et al., 2020; Geiger et al., 2020) or activation patching Meng et al. (2022), have shown promising results in uncovering causal mechanisms of LMs. They play important roles in recent interpretability research of LMs such as causal abstraction Geiger et al. (2021, 2023b), causal tracing to locate factual knowledge Meng et al. (2022); Geva et al. (2023), path patching or causal scrubbing to find causal circuits Chan et al. (2022); Conmy et al. (2023); Goldowsky-Dill et al. (2023) and distributed alignment search Geiger et al. (2023b). Previous works suggest that activation interventions that result in systematic counterfactual behaviors provide clear causal insights into model components.
@@
LMs learn highly distributed representations that encode multiple concepts in a direction Smolensky (1988); Olah et al. (2020); Elhage et al. (2022). Various methods have been proposed to find components that capture a concept, such as finding a linear subspace that modifies a concept Ravfogel et al. (2020, 2022); Belrose et al. (2023); Cao et al. (2020); Geiger et al. (2023b) and generating a sparse feature space where each direction captures a word sense or is more interpretable Arora et al. (2018); Bricken et al. (2023); Cunningham et al. (2024); Tamkin et al. (2023). However, these methods have not been evaluated against each other on their ability to isolate concepts. Isolating an individual concept is also related to the goal of “disentanglement” in representation learning Schölkopf et al. (2021), where each direction captures a single generative factor. In this work, we focus on isolating the causal effect of a representation.
@@ Appendix A Dataset Details
Attributes
$|A_{E}|$
Sample Values
Sample Prompts
@@
Country
158
United States, China, Russia, Brazil, Australia
city to country: Toronto is in Canada. {E} is in, [{"city": "Paris", "country": "France"}, {"city": "{E}",   "country": "
Continent
6
Asia, Europe, Africa, North America, South America
{E} is a city in the continent of, [{"city": "{E}", "continent": "
Latitude
122
41, 37, 47, 36, 35
[{"city": "Rio de Janeiro", "lat": "23"},   {"city": "{E}", "lat": ", [{"city": "{E}", "lat": "
Longitude
317
30, 9, 10, 33, 11
[{"city": "Rome", "long": "12.5"}, {"city": "{E}", "long": ", "long": "122.4"}, {"city": "{E}", "long": "
Language
159
English, Spanish, Chinese, Russian, Portuguese
[{"city": "Beijing", "lang": "Chinese"},   {"city": "{E}", "lang": ", [{"city": "{E}", "official language": "
Timezone
267
America/Chicago, Asia/Shanghai, Asia/Kolkata, Europe/Moscow, America/Sao_Paulo
Time zone in Los Angeles is America/Santiago;   Time zone in {E} is, [{"city": "New Delhi", "timezone": "UTC+5:30"},   {"city": "{E}", "timezone": "UTC
@@
Field
7
Medicine, Physics, Chemistry, Literature, Peace
Jules A. Hoffmann won the Nobel Prize in Medicine.   {E} won the Nobel Prize in, name: {E}, award: Nobel Prize in
Award Year
118
2001, 2019, 2009, 2011, 2000
"name": {E}, "award": "Nobel Prize", "year": ", laureate: Frances H. Arnold, year: 2018, laureate: {E}, year:
Birth Year
145
1918, 1940, 1943, 1911, 1941
Alan Heeger was born in 1936. {E} was born in, laureate: {E}, date of birth (YYYY-MM-DD):
Country of Birth
81
United States, United Kingdom, Germany, France, Sweden
name: {E}, country:, Roderick MacKinnon was born in United States. {E} was born in
Gender
4
his, male, female, her
name: {E}, gender:, David M. Lee: for his contributions in physics. {E}: for
@@
Attributes
$|A_{E}|$
Sample Values
Sample Prompts
@@
Definition
986
take hold of, make certain, show, express in words, make
talk: communicate by speaking; win: achieve victory; {E}:, like: have a positive preference; walk: move on foot; {E}:
Past Tense
986
expanded, sealed, terminated, escaped, answered
present tense: {E}, past tense:, write: wrote; look: looked; {E}:
Pronun- ciation
986
@@
[…]
Singular
986
compensates, kicks, hunts, earns, accompanies
tell: tells; create: creates; {E}:, present tense: {E}, 3rd person present:
@@
Category
29
plant, non-living thing, animal, NO, fruit
bird is a type of animal: YES; rock is a type of animal: NO;   {E} is a type of animal:, Among the categories "plant", "animal", and "non-living thing",   {E} belongs to "
Color
12
green, white, yellow, brown, black
The color of apple is usually red. The color of violet is   usually purple. The color of {E} is usually, The color of apple is usually red. The color of turquoise   is usually blue. The color of {E} is usually
Size
4
cm, mm, m, km
Among the units "mm", "cm", "m", and "km",   the size of {E} is usually on the scale of ", Given the units "mm" "cm" "m" and "km",   the size of {E} usually is in "
Texture
2
soft, hard
hard or soft: rock is hard; towel is soft;   blackberry is soft; wood is hard; {E} is, Texture: rock: hard; towel: soft; blackberry: soft;   charcoal: hard; {E}:
@@
Duty
650
treat patients, teach students, sell products, create art, serve food
"occupation": "photographer", "duties": "to capture   images using cameras"; "occupation": "{E}", "duties": "to, "occupation": "{E}", "primary duties": "to
Gender Bias
9
he, male, his, female, she
The {E} left early because The newspaper praised the {E} for
Industry
280
construction, automotive, education, health care, agriculture
"occupation": "sales manager", "industry": "retail";   "occupation": "{E}", "industry": ", "occupation": "software developer", "industry":   "technology"; "occupation": "{E}", "industry": "
Work Location
128
office, factory, hospital, construction site, studio
"occupation": "software developer", "environment": "office";   "occupation": "{E}", "environment": "

### v2

@@ Abstract
Mor Geva Tel Aviv University morgeva@tauex.tau.ac.il                        Atticus Geiger Pr(Ai)2R Group atticusg@gmail.com
@@ 1 Introduction
The metric we use to assess interpretability methods is based on interchange interventions (also known as activation patching). This operation has emerged as a workhorse in interpretability, with a wide swath of research applying the technique to test if a high-level concept is stored in a model representation and used during inference Geiger et al. (2020); Vig et al. (2020); Geiger et al. (2021); Li et al. (2021); Finlayson et al. (2021); Meng et al. (2022); Chan et al. (2022); Geva et al. (2023); Wang et al. (2023); Hanna et al. (2023); Conmy et al. (2023); Goldowsky-Dill et al. (2023); Hase et al. (2023); Todd et al. (2024); Feng and Steinhardt (2024); Cunningham et al. (2024); Huang et al. (2023); Tigges et al. (2023); Lieberum et al. (2023); Davies et al. (2023); Hendel et al. (2023); Ghandeharioun et al. (2024).
Specifically, we use the LM to process a prompt like “Paris is in the continent of” and then intervene on the neurons $\mathbf{N}$ to fix the feature $F$ to be the value it would have if the LM were given a prompt like “Tokyo is a large city.” If this leads the LM to output “Asia” instead of “Europe”, then we have evidence that the feature $F$ encodes the attribute “continent”. Then, we perform the same intervention when the LM processes a prompt like “People in Paris speak”. If the LM outputs “French” rather than “Japanese’, then we have evidence that the feature $F$ has disentangled the attributes “continent” and “language”.
@@ 2 The Ravel Dataset
The goal of Ravel is to assess the ability of methods to isolate individual explanatory factors in model representations (desideratum 4), and do so in a way that is faithful to how the target models work (desideratum 1). The dataset train/test structure seeks to ensure that methods are evaluated by how well their explanations generalize to new cases (desideratum 3), and Ravel is designed to support intervention-based metrics that assess the extent to which methods have found representations that causally affect the model behavior (desideratum 2).
Entity Type Attributes # Entities # Prompt Templates
City Country, Language, Latitude, Longitude, Timezone, Continent 3552 150
Nobel Laureate Award Year, Birth Year, Country of Birth, Field, Gender 928 100
Verb Definition, Past Tense, Pronunciation, Singular 986 60
Physical Object Biological Category, Color, Size, Texture 563 60
Occupation Duty, Gender Bias, Industry, Work Location 799 50
@@ 2.1 Data Generation
Ravel offers two settings, Entity and Context, to evaluate the generalizability (desideratum 3) of an interpretability method across unseen entities and contexts. Each setting has a predefined train/dev/test structure. In Entity, for each entity type, we randomly split the entities into 50%/25%/25% for train/dev/test, but use the same set of prompt templates across the three splits. In Context, for each attribute, we randomly split the prompt templates into 50%/25%/25%, but use the same set of entities across the three splits.
@@ 2.2 Interpretability Evaluation
A central goal of Ravel is to assess methods by the extent to which they provide causal explanations of model behaviors (desideratum 2). To build such analyses, we need to put models into counterfactual states that allow us to isolate the causal effects of interest.
@@
Geiger et al. (2023b) generalize this operation to intervene upon features that are distributed across neurons using a bijective featurizer $\mathcal{F}$. Let $\mathcal{M}_{F\leftarrow f}$ be the model where neurons $\mathbf{N}$ are projected into a feature space using $\mathcal{F}$, the feature $F$ is fixed to take on value $f$, and then the features are projected back into the space of neural activations using $\mathcal{F}^{-1}$. If we let $\tau(\mathcal{M}(x))$ be the token that a model predicts for a given prompt $x\in\mathcal{D}{}{}$, then comparisons between $\tau(\mathcal{M}(x))$ and $\tau(\mathcal{M}_{F\leftarrow f}(x))$ yield insights into the causal role that $F$ plays in model behavior.
@@
If $F$ encodes $A$, then interventions on $F$ should change the value of $A$. When $\mathcal{M}$ is given a prompt $x\in\mathcal{P}_{E}^{A}$, we can intervene on $F$ to set the value to what it would be if a second prompt $x^{\prime}\in\mathcal{P}_{E^{\prime}}^{A^{\prime}}\cup\mathcal{W}_{E^{\prime}}$ were provided. The token predicted by $\mathcal{M}$ should change from $A_{E}$ to $A_{E^{\prime}}$:
@@
To balance these two objectives, we define the Disentangle score as a weighted average between Cause and Iso.
@@
In practice, two attributes might not be fully disentanglable in the model $\mathcal{M}$ so there is no guarantee that it is possible to find a feature $F$ that achieves $\texttt{Cause}=1$ and $\texttt{Iso}=1$ at the same time. However, evidence that two attributes might not be separable is an insight into how knowledge is structured in the model.
@@ 3.2 Sparse Autoencoder
$\sum_{x\in\mathcal{D}}||\texttt{GetVals}(\mathcal{M}(x),\mathbf{N})-\Big{(}W_{2}\mathbf{f}+b_{2}\Big{)}||_{2}+||\mathbf{f}||_{1}$
$\mathbf{f}=\mathsf{ReLU}(W_{1}(\texttt{GetVals}(\mathcal{M}(x),\mathbf{N})-b_{2})+b_{1})$
with $W_{1}\in\mathbb{R}^{n\times m}$, $W_{2}\in\mathbb{R}^{m\times n}$, $b_{1}\in\mathbb{R}^{m}$, and $b_{2}\in\mathbb{R}^{n}$. To construct a training datset, we sample 100k sentences from the Wikipedia corpus for each entity type, each containing a mention of an entity in the training set. We extract the 4096-dimension hidden states of Llama2-7B at the target intervention site as the input for training a sparse autoencoder with 16384 features.
We use the autoencoder to define a featurizer
$\mathcal{F}_{A}(\mathbf{n})=\mathsf{ReLU}(W_{1}(\mathbf{n}-b_{2})+b_{1})$
@@ 3.3 Relaxed Linear Adversarial Probe
We define the $\mathcal{F}$ using the set of $k$ orthonormal vectors that span the row space of $W$ and the set of $n-k$ orthonormal vectors that span the null space:
@@
Our feature $F_{A}$ is the first $k$ dimensions of the feature space, i.e. the row space of $W$. Intuitively, since the linear probe was trained to extract the attribute $A$, the rowspace is the linear subspace of neural activations that the probe is “looking at” to make predictions.
@@ 3.5 Distributed Alignment Search
$\mathcal{L}_{\texttt{Cause}}(A,F_{A},\mathcal{M})=\texttt{CE}(\texttt{II}(\mathcal{M},F_{A},x,x^{\prime}),A_{E^{\prime}})$
@@ 4.2 Results
We evaluate each method on every representation $\mathbf{N}^{L}_{t_{E}}$ and report the highest disentanglement score on test splits in Table 2. We additionally include a baseline that simply replaces the full representation $\mathbf{N}^{L}_{t_{E}}$ regardless of what attribute is being targetted (see Full Rep. in Table 2). A breakdown of the results with per-attribute Cause and Iso is in Appendix C.
@@
MDAS is the state-of-the-art method on Ravel, being able to achieve high Disentangle scores while only intervening on a feature $F_{A}$ with a dimensionality that is $4\%$ of $|\mathbf{N}|$ where $\mathbf{N}$ are the neurons the feature is distributed across (Figure 2). DBM, MDBM, and DAS, the other methods that are trained with interventions using counterfactual labels as supervision, achieve the next best performance. PCA and Sparse Autoencoder achieve the lowest Disentangle scores, which aligns with the prior finding that disentangled representations are difficult to learn without supervision Locatello et al. (2018). Unsurprisingly, more supervision results in higher performance.
@@ 5 Related Work
Intervention-based techniques, branching off from interchange intervention (Vig et al., 2020; Geiger et al., 2020) or activation patching Meng et al. (2022), have shown promising results in uncovering causal mechanisms of LMs. They play important roles in recent interpretability research of LMs such as causal abstraction Geiger et al. (2021, 2023b), causal tracing to locate factual knowledge Meng et al. (2022); Geva et al. (2023), path patching or causal scrubbing to find causal circuits Chan et al. (2022); Conmy et al. (2023); Goldowsky-Dill et al. (2023), and Distributed Alignment Search Geiger et al. (2023b). Previous works suggest that activation interventions that result in systematic counterfactual behaviors provide clear causal insights into model components.
@@
LMs learn highly distributed representations that encode multiple concepts in a overlapping sets of neurons Smolensky (1988); Olah et al. (2020); Elhage et al. (2022). Various methods have been proposed to find components that capture a concept, such as finding a linear subspace that modifies a concept Ravfogel et al. (2020, 2022); Belrose et al. (2023); Cao et al. (2020); Geiger et al. (2023b) and generating a sparse feature space where each direction captures a word sense or is more interpretable Arora et al. (2018); Bricken et al. (2023); Cunningham et al. (2024); Tamkin et al. (2023). However, these methods have not been evaluated against each other on their ability to isolate concepts. Isolating an individual concept is also related to the goal of “disentanglement” in representation learning Schölkopf et al. (2021), where each direction captures a single generative factor. In this work, we focus on isolating the causal effect of a representation.
@@ Appendix A Dataset Details
Attributes $|A_{E}|$ Sample Values Sample Prompts
@@
Country 158 United States, China, Russia, Brazil, Australia city to country: Toronto is in Canada. {E} is in, [{"city": "Paris", "country": "France"}, {"city": "{E}",   "country": "
Continent 6 Asia, Europe, Africa, North America, South America {E} is a city in the continent of, [{"city": "{E}", "continent": "
Latitude 122 41, 37, 47, 36, 35 [{"city": "Rio de Janeiro", "lat": "23"},   {"city": "{E}", "lat": ", [{"city": "{E}", "lat": "
Longitude 317 30, 9, 10, 33, 11 [{"city": "Rome", "long": "12.5"}, {"city": "{E}", "long": ", "long": "122.4"}, {"city": "{E}", "long": "
Language 159 English, Spanish, Chinese, Russian, Portuguese [{"city": "Beijing", "lang": "Chinese"},   {"city": "{E}", "lang": ", [{"city": "{E}", "official language": "
Timezone 267 America/Chicago, Asia/Shanghai, Asia/Kolkata, Europe/Moscow, America/Sao_Paulo Time zone in Los Angeles is America/Santiago;   Time zone in {E} is, [{"city": "New Delhi", "timezone": "UTC+5:30"},   {"city": "{E}", "timezone": "UTC
@@
Field 7 Medicine, Physics, Chemistry, Literature, Peace Jules A. Hoffmann won the Nobel Prize in Medicine.   {E} won the Nobel Prize in, name: {E}, award: Nobel Prize in
Award Year 118 2001, 2019, 2009, 2011, 2000 "name": {E}, "award": "Nobel Prize", "year": ", laureate: Frances H. Arnold, year: 2018, laureate: {E}, year:
Birth Year 145 1918, 1940, 1943, 1911, 1941 Alan Heeger was born in 1936. {E} was born in, laureate: {E}, date of birth (YYYY-MM-DD):
Country of Birth 81 United States, United Kingdom, Germany, France, Sweden name: {E}, country:, Roderick MacKinnon was born in United States. {E} was born in
Gender 4 his, male, female, her name: {E}, gender:, David M. Lee: for his contributions in physics. {E}: for
@@
Attributes $|A_{E}|$ Sample Values Sample Prompts
@@
Definition 986 take hold of, make certain, show, express in words, make talk: communicate by speaking; win: achieve victory; {E}:, like: have a positive preference; walk: move on foot; {E}:
Past Tense 986 expanded, sealed, terminated, escaped, answered present tense: {E}, past tense:, write: wrote; look: looked; {E}:
@@
Singular 986 compensates, kicks, hunts, earns, accompanies tell: tells; create: creates; {E}:, present tense: {E}, 3rd person present:
@@
Category 29 plant, non-living thing, animal, NO, fruit bird is a type of animal: YES; rock is a type of animal: NO;   {E} is a type of animal:, Among the categories "plant", "animal", and "non-living thing",   {E} belongs to "
Color 12 green, white, yellow, brown, black The color of apple is usually red. The color of violet is   usually purple. The color of {E} is usually, The color of apple is usually red. The color of turquoise   is usually blue. The color of {E} is usually
Size 4 cm, mm, m, km Among the units "mm", "cm", "m", and "km",   the size of {E} is usually on the scale of ", Given the units "mm" "cm" "m" and "km",   the size of {E} usually is in "
Texture 2 soft, hard hard or soft: rock is hard; towel is soft;   blackberry is soft; wood is hard; {E} is, Texture: rock: hard; towel: soft; blackberry: soft;   charcoal: hard; {E}:
@@
Duty 650 treat patients, teach students, sell products, create art, serve food "occupation": "photographer", "duties": "to capture   images using cameras"; "occupation": "{E}", "duties": "to, "occupation": "{E}", "primary duties": "to
Gender Bias 9 he, male, his, female, she The {E} left early because The newspaper praised the {E} for
Industry 280 construction, automotive, education, health care, agriculture "occupation": "sales manager", "industry": "retail";   "occupation": "{E}", "industry": ", "occupation": "software developer", "industry":   "technology"; "occupation": "{E}", "industry": "
Work Location 128 office, factory, hospital, construction site, studio "occupation": "software developer", "environment": "office";   "occupation": "{E}", "environment": "